Grade 10Physical Chemistry

B10 and B 11 are two isotopes of boron if average mass is 10.2 what is percentage of each

Rajat
9 Years agoGrade 10

1 Answer

Abhijith Sajeev
9 Years ago
Let he percentage existence of B10 isotope be x . Therefore percentage existence of the other isotope is (100-x)
 
mass of B10 isotope is 10amu
mass of B11 isotope is 11 amu
 
therefore 
10x/100+11(100-x)/100=10.2
 
10x+11(100-x)=1020
1100-x=1020
 
 
x=80
 
therefore percentage of Boron10 isotope is 80%
and Boron 11 isotope is(100-x)=20%
 
Thankyou